(-8 p^2q)(-4p^4q^5)=?
Answer by AnlytcPhil(1116)

(-8p^2q)(-4p^4q^5)=?

(-8p2q)(-4p4q5)

Put each factor in separate parentheses:

(-8)(p2)(q)(-4)(p4)(q5)

Rearrange them with numbers first, p's next, q's last

(-8)(-4)(p2)(p4)(q)(q5)

Multiply the numbers (-8)(-4) = +32

32(p2)(p4)(q)(q5)

To multiply (p2)(p4), add the exponents p2+4 = p6

32p6(q)(q5)

Give the q the exponent 1

32p6(q1)(q5)

To multiply (q1)(q5), add the exponents q1+5 = q6

32p6q6

Edwin
AnlytcPhil@aol.com